OpenStudy (umulas):

Can someone check if this is correct? Analyze the function f(x) = - tan 4x. Include: - Domain and range - Period - Two Vertical Asymptotes Alright so I first know that Period is: pi/b = pi/4

OpenStudy (umulas):

I've seen the mistake for VA I believe. retricemptote is: pi/2 quotient 4 = pi/8 3pi/2 quotient 4 = 3pi/8 Can anyone help me with this?

OpenStudy (umulas):

I meant to say asyymptote

OpenStudy (umulas):

Hmmm; I believe I can just put Range: All real numbers Domain: All Real numbers except multiple of pi/2?

zepdrix (zepdrix):

Hmm I think you were on the right track before, The asymptotes are at pi/8, 3pi/8, 5pi/8, ... So we could simplify their location by saying, \(\large \dfrac{\pi}{8}\pm\dfrac{\pi}{2}k\)

zepdrix (zepdrix):

So the domain would be .. yah the way you were thinking, all reals except .. those asymptotes, the fancy way we wrote it out.

zepdrix (zepdrix):

It might look a little nicer if we wrote it in set notation, but no big deal, \[\large D=\left\{ x \;|\; x\ne \frac{\pi}{8}\pm\frac{\pi}{2}k \right\}\] This translates to, "The Domain D is the set of all points x, such that x does not equal ..that "

OpenStudy (umulas):

Awesome! I'm going to give you the best responce, but I just have one question (forgive me, my teacher teaches me ABSOLUTELY nothing). What would I put for the D then, Domain is all numbers except multitude of pi/2, is that correct?

zepdrix (zepdrix):

Bahh I did that wrong. My bad. There should be one asymptote popping up every `period`. You correctly identified asymptotes at \(\large \dfrac{\pi}{8}\) and \(\large \dfrac{3\pi}{8}\). Notice that those are \(\large \dfrac{\pi}{4}\) away from each other? A new asymptote will pop up every \(\large \dfrac{\pi}{4}\), which is our period. It's kind of hard to say it the way you're trying to, in words, We're starting at pi/8, and then adding or subtracting any multiples of pi/4 to reach the other asymptotes. The Domain is All Reals except \(\large \dfrac{\pi}{8}\pm\dfrac{\pi}{4}k\). The k is any integer value. This ^ ugly expression is the location of each asymptote. The domain is all reals except those locations. I'm sorry, I can't think of an easier way to explain that lol.